# A Test of Substellar Evolutionary Models with High-Precision Ages from Asteroseismology and Gyrochronology for the Benchmark System HR 7672AB

### v1

@@ Abstract
1 Introduction
Brown dwarfs are objects with masses between the heaviest gas-giant planets and the lowest-mass stars ($\approx$13–75 $\text{M}_{\text{Jup}}$; e.g., Chabrier2000; Dupuy2017). They are not massive enough to sustain stable hydrogen fusion, which is the process that powers main-sequence stars. Instead, brown dwarfs can briefly fuse deuterium, and those above $\approx$65 $\text{M}_{\text{Jup}}$ can fuse lithium early in their evolution. Without a long-lived fusion source, brown dwarfs cool and fade as they radiate away their residual heat. Their effective temperature and luminosity steadily decrease with age, causing them to evolve through later spectral types (from late-M to L to T and Y). This contrasts with main-sequence stars, which maintain stable luminosities for millions to trillions of years.
The cooling of brown dwarfs is predicted by evolutionary models. Early models used simplified treatments of interior and atmospheric physics, but the low temperatures of brown dwarfs require more sophisticated approaches. Modern models incorporate non-grey atmospheres (e.g., Burrows1997; Chabrier1997; 2021ApJ...920...85M), dust and cloud physics (e.g., Chabrier2000; Baraffe2002; Saumon2008; 2012ApJ...756..172M), disequilibrium chemistry (e.g., 2006ApJ...647..552S; 2020A&A...637A..38P), vertical mixing (e.g., 2015ApJ...804L..17T), detailed equations of state and degeneracy (e.g., Chabrier2023), and assumptions about initial conditions (e.g. hot vs. cold start; e.g., 2007ApJ...655..541M).
@@ 1 Introduction
Given the wide-spread use of such models, testing their predictions is imperative. To do so, model-independent measurements of brown dwarf properties are needed, in particular the 3 key quantities of luminosity, age, and mass. Systems with at least 2 of these quantities measured are often referred to as ”benchmarks” (e.g. liu08-2m1534orbit; Zhangzj2020). Luminosities are relatively easy to measure using photometry and/or spectroscopy, supplemented by bolometric corrections or fitting model atmospheres (e.g. 2023ApJ...959...63S). However, age and mass are not directly observable for most brown dwarfs, which are free-floating objects. For a minority fraction of known objects, ages can be established via their physical association with a star or stellar association (e.g. 2006MNRAS.368.1281P). And for an even smaller fraction of the known objects, dynamical masses can be measured via the visual/astrometric orbits of binary brown dwarfs (e.g., Dupuy2017) and brown dwarfs that are companions to nearby stars (e.g. Brandt2019).
@@
More recently, HR 7672B’s atmosphere has been studied with AO-assisted high-spectral resolution measurements. Wangj2022 used $K$-band spectroscopy with Keck/KPIC to measure carbon and oxygen abundances, key to CO, H2O, and CH4 chemistry. The elemental abundances of the host star agree well with the host star, supporting a star-like formation scenario for the companion. Kasagi2025 further obtained high-resolution $YJH$-band spectra with Subaru/REACH. Their atmospheric retrievals detected H2O and FeH absorption and required an optically thick cloud layer to reproduce the data, consistent with L-dwarf cloud physics. They inferred cloud-top temperatures suggestive of condensates such as TiO2, Al2O3, or Fe.
In this study, we derive the age of HR 7672A — and hence the system — using asteroseismology from three consecutive nights of Keck Planet Finder (KPF) radial velocity data and with gyrochronology based on stellar rotation (§2–§3). We also refine the dynamical mass of HR 7672B, using a new epoch of AO imaging combined with additional RV data for the primary. These new high-precision measurements of the system enable strong tests of current substellar evolutionary models (§4). We also discuss the nearly-simultaneous KPF and TESS observations, which offer additional insights into stellar pulsations (§5).
2 Observations
@@
2.1 Asteroseismology with Keck Planet Finder
@@
Asteroseismology is the study of stellar oscillations, which probe internal structure and provide precise fundamental stellar properties (Aerts2010; Basu2017). Extremely Precise Radial Velocity (EPRV) techniques now achieve radial velocity measurements with precisions down to 30 cm/s for individual measurements (Fischer2016; Wright2018). At this precision, stars that oscillate at amplitudes too low to be detected from photometry like Kepler (Gilliland2011) and TESS (Ricker2015), such as main-sequence K dwarfs, can now be routinely detected via the EPRV method. Oscillations have recently been detected in K-type main-sequence dwarfs with amplitudes of only 1–6 cm/s in $\epsilon$ Ind A (Campante2024; Lundkvist2024), $\sigma$ Dra (Hon2024-sig-dra) and HD 219134 (Liyg2025), as well as in the subgiant HD 118203 (Zhangjw2024).
@@
The advantage of the EPRV method is particularly great for young, magnetically active stars (Chaplin2011-B; Mathur2019). Their oscillation amplitudes are, in general, smaller than those of older stars (Huber2011; Campante2014; Sayeed2025), so they can often be obscured by granulation or white noise in photometry (Sulis2023). EPRV is less sensitive to granulation at oscillation frequencies of interest, enabling higher signal-to-noise ratio (SNR) of oscillations in such targets. The star we are studying, HR 7672A, is a moderately young G-type dwarf that illustrates this capability.
@@
We obtained approximately 8 hours of observations per night over 3 consecutive nights, from July 12 to 14, 2024 (Hawaii Standard Time), using the Keck Planet Finder (KPF) mounted on the Keck I telescope (gibson2016-kpf; gibson2018-kpf; gibson2020-kpf; gibson2024-kpf). A total of 1,084 radial velocity measurements were collected. Observing conditions were moderately good, but with occasional gusts resulting in the poorest seeing reaching $2.0^{\prime\prime}$. To maintain the signal-to-noise ratio to reach adequate RV precision per spectra under the variable conditions, we used a mixture of 60-second and 90-second exposures, each followed by a 15-second readout, yielding cadences of 75 or 105 seconds, where the latter corresponds to a Nyquist frequency of 4800 $\mu{\rm Hz}$. The average time-domain scatter measured using the power density at high frequency is 0.68 m/s. Table 1 lists these high-cadence radial velocity measurements obtained with the Keck Planet Finder.
@@
Figure 1 shows the RV time series, along with the power spectrum calculated using the Lomb-Scargle method (lomb1976; scargle1982), using the reported RV uncertainties as weights (Frandsen1995). We see a clear power excess around 3000 $\mu{\rm Hz}$. Based on a heavily smoothed power spectrum (Kjeldsen2008), we measured the frequency of maximum power as ${\nu_{\text{max}}}=3245\pm 75$ $\mu{\rm Hz}$. We note that this measured $\nu_{\text{max}}$ does not translate to stellar properties accurately, because the value of $\nu_{\text{max}}$ depends on the technique (photometric bandpass or selected spectroscopic lines; Hon2024-sig-dra), stellar-cycle variability (Howe2020), and time variability introduced by stochastic excitation (Jiangc2023; Sreenivas2024).
@@
The oscillations observed around $\nu_{\text{max}}$ are acoustic modes, and their frequencies can be well described by the asymptotic relation involving two quantum numbers, the radial order, $n$, and angular degree, $\ell$ (Tassoul1980; Scherrer1983):
@@
The large frequency separation, $\Delta\nu$, depends primarily on the sound travel time across the star and, to first order, scales with the square root of the mean stellar density (Ulrich1986). The phase offset, $\epsilon$, generally falls between 0.8 and 1.6 in main-sequence stars (White2012), and depends on the turning points of modes (Roxburgh2003). The small separations, $\delta\nu_{0,\ell}$, quantify the frequency offsets between modes of different $\ell$ but with adjacent $n$. In main-sequence stars, $\delta\nu_{0,\ell}$ is sensitive to the chemical gradient near the core (Tassoul1994; Roxburgh1994), making it a useful diagnostic for tracing hydrogen-burning processes and, therefore, for determining stellar ages on the main sequence (jcd1984-02; White2011; Bellinger2019; Hon2024-flow).
@@
To identify and measure these oscillation modes, we followed the method outlined in Liyg2025. We applied Gold deconvolution to the power spectrum with the spectral window to reveal regular structures of mode spacings (gold1964; morhac2003-gold; Ong et al. in preparation). Figure 2(a) shows the power spectrum displayed in an échelle format, which involves folding into segments of $\Delta\nu$ and stacking them vertically (the échelle is duplicated for clarity). Figure 2(b) shows the deconvolved spectrum, revealing clear $\ell=0$33 mode ridges, corresponding to the regular spacings dictated by Equation 1 (see also Figure 1(c)). Figure 2(c) shows the collapsed power spectrum from summing power along the vertical direction in Figure 2(b). Because the small spacings $\delta\nu_{02}$ in main-sequence stars are close to 1 c/d, which is the spacing between sidelobes (see inset of Figure 10), the sidelobes of the $l=0$ modes overlap with the main signal of the $l=2$ modes, and vice versa. Therefore it is very difficult to identify the true oscillation frequency from the original power spectrum (Figure 2a). The Gold deconvolution provides a clean way to suppress power from the sidelobes and thus reveal the true mode structure.
@@
From Figures 2(b) and (c), we identified the frequencies of peaks along each ridge, guided by the asymptotic relation (Equation 1), which predicts that modes of the same $\ell$ degree should align vertically. We then constructed a time-series model consisting of multiple sine waves, with frequencies fixed to those identified from the deconvolved spectrum and amplitudes $a$ and phases $\phi$ as free parameters. We fitted this model to the RV time series by minimizing RV-uncertainty-weighted $\chi^{2}$ statistic. We estimated the average noise level $\sigma_{\rm noise}$ to be 0.03 m/s, based on the amplitude spectrum at high frequencies (>5000>5000 $\mu{\rm Hz}$). Modes with amplitudes exceeding 3.5 times $\sigma_{\rm noise}$ were retained in the final list of modes. These are highlighted in Figure 2(d). We then computed the power spectrum of the RV time series after subtracting the best-fitting model. The corresponding deconvolved spectrum is shown in Figure 2(e), and shows no clear remaining structure. The lack of any clear residual structure in Figure 2(e) indicates that treating the modes as coherent over the observing window is adequate. We estimated the uncertainties of the frequencies using an analytical expression (Montgomery1999; Kjeldsen2012):
@@
where $T_{\rm obs}$ is the effective observing duration and $\tau$ is the mode lifetime, which we assume to be 3 d, based on its dependency on effective temperature (Chaplin2009). Although the formal frequency resolution is set by $1/T_{\rm obs}$, the uncertainty of an extracted mode frequency can be much smaller. The resolution specifies the minimum spacing at which two unresolved peaks can be distinguished in the power spectrum, but it does not limit the precision with which the centroid of a single peak can be measured. We also corrected the line-of-sight Doppler shifts for these frequencies following davies2014-los, using a radial velocity of $5.6685$ km/s from an average value over three nights. The final list of frequencies are given in Table 2. Fitting Equation 1 to these frequencies yields the following asymptotic parameters: ${\Delta\nu}{}=132.798\pm 0.122$ $\mu{\rm Hz}$, $\epsilon=1.336\pm 0.022$, ${\delta\nu_{0,1}}=5.43\pm 0.69$ $\mu{\rm Hz}$, ${\delta\nu_{0,2}}=12.45\pm 0.64$ $\mu{\rm Hz}$, and ${\delta\nu_{0,3}}=22.80\pm 1.16$ $\mu{\rm Hz}$.
2.2 Dynamical Orbit with NIRC2 and Radial Velocities

### v2

@@ Abstract
I Introduction
Brown dwarfs are objects with masses between the heaviest gas-giant planets and the lowest-mass stars ($\approx$13–75 $\text{M}_{\text{Jup}}$; e.g., G. Chabrier & I. Baraffe, 2000; T. J. Dupuy & M. C. Liu, 2017a). They are not massive enough to sustain stable hydrogen fusion, which is the process that powers main-sequence stars. Instead, brown dwarfs can briefly fuse deuterium, and those above $\approx$65 $\text{M}_{\text{Jup}}$ can fuse lithium early in their evolution. Without a long-lived fusion source, brown dwarfs cool and fade as they radiate away their residual heat. Their effective temperature and luminosity steadily decrease with age, causing them to evolve through later spectral types (from late-M to L to T and Y). This contrasts with main-sequence stars, which maintain stable luminosities for millions to trillions of years.
The cooling of brown dwarfs is predicted by evolutionary models. Early models used simplified treatments of interior and atmospheric physics, but the low temperatures of brown dwarfs require more sophisticated approaches. Modern models incorporate non-grey atmospheres (e.g., A. Burrows et al., 1997; G. Chabrier & I. Baraffe, 1997; M. S. Marley et al., 2021), dust and cloud physics (e.g., G. Chabrier & I. Baraffe, 2000; I. Baraffe et al., 2002; D. Saumon & M. S. Marley, 2008; C. V. Morley et al., 2012), disequilibrium chemistry (e.g., D. Saumon et al., 2006; M. W. Phillips et al., 2020a), vertical mixing (e.g., P. Tremblin et al., 2015), detailed equations of state and degeneracy (e.g., G. Chabrier et al., 2023), and assumptions about initial conditions (e.g. hot vs. cold start; e.g., M. S. Marley et al., 2007).
@@ I Introduction
Given the wide-spread use of such models, testing their predictions is imperative. To do so, model-independent measurements of brown dwarf properties are needed, in particular the 3 key quantities of luminosity, age, and mass. Systems with at least 2 of these quantities measured are often referred to as “benchmarks” (e.g. M. C. Liu et al., 2008; Z. Zhang et al., 2020). Luminosities are relatively easy to measure using photometry and/or spectroscopy, supplemented by bolometric corrections or fitting model atmospheres (e.g. A. Sanghi et al., 2023). However, age and mass are not directly observable for most brown dwarfs, which are free-floating objects. For a minority fraction of known objects, ages can be established via their physical association with a star or stellar association (e.g. D. J. Pinfield et al., 2006). And for an even smaller fraction of the known objects, dynamical masses can be measured via the visual/astrometric orbits of binary brown dwarfs (e.g., T. J. Dupuy & M. C. Liu, 2017a) and brown dwarfs that are companions to nearby stars (e.g. T. D. Brandt et al., 2019a).
@@
More recently, HR 7672B’s atmosphere has been studied with AO-assisted high-spectral resolution measurements. J. Wang et al. (2022) used $K$-band spectroscopy with Keck/KPIC to measure carbon and oxygen abundances, key to CO, H2O, and CH4 chemistry. The elemental abundances of the host star agree well with the host star, supporting a star-like formation scenario for the companion. Y. Kasagi et al. (2025) further obtained high-resolution $YJH$-band spectra with Subaru/REACH. Their atmospheric retrievals detected H2O and FeH absorption and required an optically thick cloud layer to reproduce the data, consistent with L-dwarf cloud physics. They inferred cloud-top temperatures suggestive of condensates such as TiO2, Al2O3, or Fe.
In this study, we derive the age of HR 7672A — and hence the system — using asteroseismology from three consecutive nights of Keck Planet Finder (KPF) radial velocity data and with gyrochronology based on stellar rotation (§II–§III). We also refine the dynamical mass of HR 7672B, using a new epoch of AO imaging combined with additional RV data for the primary. These new high-precision measurements of the system enable strong tests of current substellar evolutionary models (§IV). We also discuss the nearly-simultaneous KPF and TESS observations, which offer additional insights into stellar pulsations (§II.2).
II Observations
@@
II.1 Asteroseismology with Keck Planet Finder
@@
Asteroseismology is the study of stellar oscillations, which probe internal structure and provide precise fundamental stellar properties (C. Aerts et al., 2010; S. Basu & W. J. Chaplin, 2017). Extremely Precise Radial Velocity (EPRV) techniques now achieve radial velocity measurements with precisions down to 30 cm/s for individual measurements (D. A. Fischer et al., 2016; J. T. Wright, 2018). At this precision, stars that oscillate at amplitudes too low to be detected from photometry like Kepler (R. L. Gilliland et al., 2011) and TESS (G. R. Ricker et al., 2015), such as main-sequence K dwarfs, can now be routinely detected via the EPRV method. Oscillations have recently been detected in K-type main-sequence dwarfs with amplitudes of only 1–6 cm/s in $\epsilon$ Ind A (T. L. Campante et al., 2024; M. S. Lundkvist et al., 2024), $\sigma$ Dra (M. Hon et al., 2024b) and HD 219134 (Y. Li et al., 2025), as well as in the subgiant HD 118203 (J. Zhang et al., 2024).
@@
The advantage of the EPRV method is particularly great for young, magnetically active stars (W. J. Chaplin et al., 2011; S. Mathur et al., 2019). Their oscillation amplitudes are, in general, smaller than those of older stars (D. Huber et al., 2011a; T. L. Campante et al., 2014; M. Sayeed et al., 2025), so they can often be obscured by granulation or white noise in photometry (S. Sulis et al., 2023). EPRV is less sensitive to granulation at oscillation frequencies of interest, enabling higher signal-to-noise ratio (SNR) of oscillations in such targets. The star we are studying, HR 7672A, is a moderately young G-type dwarf that illustrates this capability.
@@
We obtained approximately 8 hours of observations per night over 3 consecutive nights, from July 12 to 14, 2024 (Hawaii Standard Time), using the Keck Planet Finder (KPF) mounted on the Keck I telescope (S. R. Gibson, 2016; S. R. Gibson et al., 2018, 2020, 2024). A total of 1,084 radial velocity measurements were collected. Observing conditions were moderately good, but with occasional gusts resulting in the poorest seeing reaching $2.0^{\prime\prime}$. To maintain the signal-to-noise ratio to reach adequate RV precision per spectra under the variable conditions, we used a mixture of 60-second and 90-second exposures, each followed by a 15-second readout, yielding cadences of 75 or 105 seconds, where the latter corresponds to a Nyquist frequency of 4800 $\mu{\rm Hz}$. The average time-domain scatter measured using the power density at high frequency is 0.68 m/s. Table 1 lists these high-cadence radial velocity measurements obtained with the Keck Planet Finder.
@@
Figure 1 shows the RV time series, along with the power spectrum calculated using the Lomb-Scargle method (N. R. Lomb, 1976; J. D. Scargle, 1982), using the reported RV uncertainties as weights (S. Frandsen et al., 1995). We see a clear power excess around 3000 $\mu{\rm Hz}$. Based on a heavily smoothed power spectrum (H. Kjeldsen et al., 2008), we measured the frequency of maximum power as ${\nu_{\text{max}}}=3245\pm 75$ $\mu{\rm Hz}$. We note that this measured $\nu_{\text{max}}$ does not translate to stellar properties accurately, because the value of $\nu_{\text{max}}$ depends on the technique (photometric bandpass or selected spectroscopic lines; M. Hon et al. 2024b), stellar-cycle variability (R. Howe et al., 2020), and time variability introduced by stochastic excitation (C. Jiang et al., 2023; K. R. Sreenivas et al., 2024).
@@
The oscillations observed around $\nu_{\text{max}}$ are acoustic modes, and their frequencies can be well described by the asymptotic relation involving two quantum numbers, the radial order, $n$, and angular degree, $\ell$ (M. Tassoul, 1980; P. H. Scherrer et al., 1983):
@@
The large frequency separation, $\Delta\nu$, depends primarily on the sound travel time across the star and, to first order, scales with the square root of the mean stellar density (R. K. Ulrich, 1986). The phase offset, $\epsilon$, generally falls between 0.8 and 1.6 in main-sequence stars (T. R. White et al., 2012), and depends on the turning points of modes (I. W. Roxburgh & S. V. Vorontsov, 2003). The small separations, $\delta\nu_{0,\ell}$, quantify the frequency offsets between modes of different $\ell$ but with adjacent $n$. In main-sequence stars, $\delta\nu_{0,\ell}$ is sensitive to the chemical gradient near the core (M. Tassoul, 1990; I. W. Roxburgh & S. V. Vorontsov, 1994), making it a useful diagnostic for tracing hydrogen-burning processes and, therefore, for determining stellar ages on the main sequence (J. Christensen-Dalsgaard, 1984; T. R. White et al., 2011; E. P. Bellinger, 2019; M. Hon et al., 2024a).
@@
To identify and measure these oscillation modes, we followed the method outlined in Y. Li et al. (2025). We applied Gold deconvolution to the power spectrum with the spectral window to reveal regular structures of mode spacings (R. Gold 1964; M. Morháč et al. 2003; Ong et al. in preparation). Figure 2(a) shows the power spectrum displayed in an échelle format, which involves folding into segments of $\Delta\nu$ and stacking them vertically (the échelle is duplicated for clarity). Figure 2(b) shows the deconvolved spectrum, revealing clear $\ell=0$33 mode ridges, corresponding to the regular spacings dictated by Equation 1 (see also Figure 1(c)). Figure 2(c) shows the collapsed power spectrum from summing power along the vertical direction in Figure 2(b). Because the small spacings $\delta\nu_{02}$ in main-sequence stars are close to 1 c/d, which is the spacing between sidelobes (see inset of Figure 3), the sidelobes of the $l=0$ modes overlap with the main signal of the $l=2$ modes, and vice versa. Therefore it is very difficult to identify the true oscillation frequency from the original power spectrum (Figure 2a). The Gold deconvolution provides a clean way to suppress power from the sidelobes and thus reveal the true mode structure.
@@
From Figures 2(b) and (c), we identified the frequencies of peaks along each ridge, guided by the asymptotic relation (Equation 1), which predicts that modes of the same $\ell$ degree should align vertically. We then constructed a time-series model consisting of multiple sine waves, with frequencies fixed to those identified from the deconvolved spectrum and amplitudes $a$ and phases $\phi$ as free parameters. We fitted this model to the RV time series by minimizing RV-uncertainty-weighted $\chi^{2}$ statistic. We estimated the average noise level $\sigma_{\rm noise}$ to be 0.03 m/s, based on the amplitude spectrum at high frequencies (>5000>5000 $\mu{\rm Hz}$). Modes with amplitudes exceeding 3.5 times $\sigma_{\rm noise}$ were retained in the final list of modes. These are highlighted in Figure 2(d). We then computed the power spectrum of the RV time series after subtracting the best-fitting model. The corresponding deconvolved spectrum is shown in Figure 2(e), and shows no clear remaining structure. The lack of any clear residual structure in Figure 2(e) indicates that treating the modes as coherent over the observing window is adequate. We estimated the uncertainties of the frequencies using an analytical expression (M. H. Montgomery & D. O’Donoghue, 1999; H. Kjeldsen & T. R. Bedding, 2012):
@@
where $T_{\rm obs}$ is the effective observing duration and $\tau$ is the mode lifetime, which we assume to be 3 d, based on its dependency on effective temperature (W. J. Chaplin et al., 2009). Although the formal frequency resolution is set by $1/T_{\rm obs}$, the uncertainty of an extracted mode frequency can be much smaller. The resolution specifies the minimum spacing at which two unresolved peaks can be distinguished in the power spectrum, but it does not limit the precision with which the centroid of a single peak can be measured.
We also corrected the line-of-sight Doppler shifts for these frequencies following G. R. Davies et al. (2014), using a radial velocity of $5.6685$ km/s from an average value over three nights. The final list of frequencies are given in Table 2. Fitting Equation 1 to these frequencies yields the following asymptotic parameters: ${\Delta\nu}{}=132.798\pm 0.122$ $\mu{\rm Hz}$, $\epsilon=1.336\pm 0.022$, ${\delta\nu_{0,1}}=5.43\pm 0.69$ $\mu{\rm Hz}$, ${\delta\nu_{0,2}}=12.45\pm 0.64$ $\mu{\rm Hz}$, and ${\delta\nu_{0,3}}=22.80\pm 1.16$ $\mu{\rm Hz}$.
II.2 Asteroseismology with TESS
Because oscillation frequencies are detected in both the TESS and KPF datasets, HR 7672 offers an opportunity to constrain photometry-to-RV amplitude ratios. These ratios have previously been measured in only a few stars with simultaneous data, including Procyon (D. Huber et al., 2011b), HD 35833 (A. F. Gupta et al., 2022), and $\beta$ Aql (H. Kjeldsen et al., 2025). We follow the approach of H. Kjeldsen et al. (2008) to measure oscillation amplitudes. Each power density spectrum — defined as the power spectrum multiplied by the effective observing duration — was smoothed with a Gaussian kernel with a full width at half-maximum of $4{\Delta\nu}{}$ to produce a single power excess hump. The smoothed spectrum was then multiplied by ${\Delta\nu}{}/c$ to estimate the power per radial order, where $c$ accounts for the effective number of modes per order (J. Ballot et al., 2011). Given the star’s similarity to the Sun, we adopt solar values calculated by H. Kjeldsen et al. (2008): $c=4.09$ for velocity data and $c=2.91$ for TESS photometry. Figure 6 shows the resulting spectra. The square root of the maximum near $\nu_{\text{max}}$ yields the oscillation amplitude. Uncertainties in the oscillation amplitudes were estimated via bootstrapping: in each realization, the original spectrum was multiplied by a noise spectrum, and the oscillation amplitude was remeasured. The spread in amplitude estimates across all realizations was adopted as the uncertainty.
For the KPF data, we measured a RV amplitude of $16.9\pm 0.6$ cm s-1. In the TESS data, oscillations are not clearly visible as a distinct power excess due to dominant contributions from granulation and white noise. We therefore derive an upper limit of $2.8\pm 0.7$ ppm for the photometric amplitude in the TESS bandpass, estimated from the maximum power around $\nu_{\text{max}}$. Applying a bolometric correction factor of 1.2 (M. N. Lund, 2019), the corresponding bolometric amplitude is $3.4\pm 0.9$ ppm. These values yield an upper limit on the bolometric photometry-to-RV amplitude ratio of $20.1\pm 7.6$ ppm (m s-1)-1. Both the amplitudes and their ratio are similar to the solar values. For comparison, the solar oscillation amplitude is approximately 18.7 cm s-1 in RV (H. Kjeldsen et al., 2008) and 3.6 ppm in photometry (E. Michel et al., 2009), corresponding to a ratio of about 19.2 ppm (m s-1)-1.
Further, both the amplitudes and their ratio are broadly in line with predictions from 3D hydrodynamic simulations (Y. Zhou et al., 2021). Repeating the methods of Y. Zhou et al. (2021), we obtained a simulated solar amplitude ratio between bolometric intensity and fluid velocity of 20.12 ppm (m s-1)-1. For a hotter main-sequence model whose basic stellar parameters are close to those of HR 7672A (${T_{\text{eff}}}{}=6000$ K and ${\log g}{}=4.5$ dex), the amplitude ratio is slightly lower at 19.27 ppm (m s-1)-1. The trend predicted from simulations is also consistent with the empirical amplitude scaling relation of H. Kjeldsen & T. R. Bedding (1995). Assuming the conversion factor between fluid velocity and RV is invariant from the Sun to HR 7672A, which is reasonable given the similarity between the two stars, the expectation is that the actual photometry-to-RV amplitude ratio of the latter is slightly smaller than 19.2 ppm (m s-1)-1, in agreement with our measured upper limit.
Oscillation amplitudes are expected to be lower in stars with higher magnetic activity (T. L. Campante et al., 2014; M. Sayeed et al., 2025). M. Sayeed et al. (2025) quantified this dependency using the chromospheric activity levels derived from Ca II H&K line emission, using the logarithmic flux ratio between the emission flux and the bolometric flux, $\log R^{\prime}_{\rm HK}$. S. Boro Saikia et al. (2018) reported $\log R^{\prime}_{\rm HK}$ values between 4.68-4.68 and 4.74-4.74 for this star; for comparison, the solar value is around 4.94-4.94 (R. Egeland et al., 2017). According to Equation 10 of M. Sayeed et al. (2025), this corresponds to an amplitude ratio between 0.81 and 0.85 relative to the Sun. Although our measurement is broadly consistent with this prediction, it does not provide conclusive evidence for a reduced amplitude in this star due to the large uncertainties. The photometric-to-RV amplitude ratio, however, is likely less affected by magnetic activities, as it depends on the stellar structure near the photosphere to a first order (Y. Zhou et al., 2021). By comparing hydrodynamics and magnetohydrodynamics simulations of stellar atmospheres, T. S. Bhatia et al. (2022) show that changes in near-surface density, temperature, and pressure stratification caused by the small-scale dynamo are less than 1% for G-type stars.
II.3 Dynamical Orbit with NIRC2 and Radial Velocities